OpenStudy (anonymous):

How would you transform P = Ae^kh into the linear form Y = mX + c? My attempt: ln(P) = ln(Ae^kh) ln(P) = kh ln(A) Here's where I got stuck, and went ln(P) = ln(A) + kh ln(P) = kh + ln(A) Which doesn't quite seem right..

OpenStudy (aravindg):

well you made a mistake in 2nd step .you cant distribute kh to front in this case you should proceed like ln(P) = ln(Ae^kh) ln(P) =ln(A)+ln(e^kh)

OpenStudy (aravindg):

now can you do it ?

OpenStudy (anonymous):

Thanks! Now: ln(P) = ln(a) + kh ln(e) And since ln(e) = 1, ln(P) = ln(A) + kh Right?

OpenStudy (aravindg):

yep :)

OpenStudy (anonymous):

So would that be the final result? Or is there a way to continue until P is on its own?

OpenStudy (aravindg):

I think that is enough you have the foem y=mx+c c=kh m=1

OpenStudy (aravindg):

*form

OpenStudy (anonymous):

I understand c = kh, but how did you determine m = 1?

hartnn (hartnn):

umm...i guess, ln(P) = ln(A) + kh y= c+ mx ln P = y, ln A =c, kh=mX now m can be taken as k or h, depends on which one is variable, the other one, which is not variable will be taken as slope, m

hartnn (hartnn):

**now X can be taken as k or h,

OpenStudy (aravindg):

actually we need info on what are the constants to determine that

OpenStudy (aravindg):

If it is given A is constant you need to consider kh as mx else if kh is constant then you need to consider log A as mx

OpenStudy (anonymous):

All the task requires is to prove that P = Ae^kh is valid for a given data set. So to prove this, we have to transform it into Y = mx + b, graph it, determine the y-intercept using the graph and use that as c. Would considering kh as mx work in this case?

hartnn (hartnn):

generally, the variable is in the power of 'e' so, yes.

OpenStudy (aravindg):

yep

OpenStudy (anonymous):

Sweet, I should be fine from here. Thanks for your help guys, I'll be back if I have any more issues :)